# Reinforcement Learning Discovers Efficient Decentralized Graph Path Search Strategies

### v1

@@ 4.2 Baselines
The learned baselines we use are the MLPA2C and MLPA2CWD techniques as introduced in Section 3.2.1. Furthermore, we consider a suite of heuristic baselines that utilize homophily and graph structure for graph path search. The simplest baseline, GreedyWalker, selects the next node greedily based on the smallest Euclidean attribute distance: $\pi(u)=\text{argmin}_{u^{\prime}\in\mathcal{N}(u)}\|\mathbf{x}_{u^{\prime}}-\mathbf{x}_{u_{\text{tgt}}}\|_{2}.$
@@ 4.3 Model Evaluation & Selection
To mitigate potential memorization of routes during training, especially when nodes are uniquely identifiable based on attributes, we partition the node set $V$ into three disjoint groups: $V^{\text{train}}$, $V^{\text{val}}$, and $V^{\text{test}}$ at ratios of $80\%/10\%/10\%$. The source node $u_{\text{src}}$ is sampled uniformly at random from $V$, while the target node $u_{\text{tgt}}$ depends on whether training, validation, or testing is performed. For training, we always sample a “fresh” source-target pair, while for validation and evaluation the source-target pairs are serialized and stored (such that the performance evaluated over them is consistent). The Mean Oracle Ratio $\bar{R}_{\text{oracle}}$ is used as the primary metric for model validation and evaluation.
@@ 5.3 Ablation of Node Representation
As shown in Table 2, GARDEN significantly outperforms the MLPA2C and MLPA2CWD designs that only use raw node attributes and MLP-parameterized policies in both $R_{\text{trunc}}$ and $\bar{R}_{\text{oracle}}$. The standard MLP-parameterized model achieves the best Win Rate $R_{\text{win}}$, and the differences with respect to this metric are less conclusive. This can be explained by the arbitrary breaking of ties performed when path lengths match for all methods, coupled with the high graph density parameter $\beta=5$ leading to shorter path lengths.
@@ A.1 Implementation Details
In a future version, we will release code that enables reproducibility of the results presented in this work. This will include instructions on how to set up the dependencies, download the publicly available data, and run the methods.
@@
A.2 Additional Tables and Figures
In Figure 3, we plot GARDEN’s learned value function across the social network ego graphs. Brighter colors indicate a higher estimated value function relative to the target node, which is indicated with a black arrow and chosen randomly from the respective test sets. For comparison, we also plot the implicit preferability score $-\|\mathbf{x}_{u}-\mathbf{x}_{u_{\text{tgt}}}\|_{2}/\tau$ generated by the best-performing baseline, DistanceWalker, for the same source-target pairs. DistanceWalker struggles with Euclidean pairwise attribute distance due to high dimensionality and sparsity of node attributes. Conversely, values obtained by GARDEN serve as a reliable proxy for graph distance, assigning highest values to nodes in the target’s cluster or clusters with strong connectivity to the target’s community.

### v2

@@ 4.2 Baselines
The learned baselines we use are the MLPA2C and MLPA2CWD techniques as introduced in Section 3.2. Furthermore, we consider a suite of heuristic baselines that utilize homophily and graph structure for graph path search. The simplest baseline, GreedyWalker, selects the next node greedily based on the smallest Euclidean attribute distance: $\pi(u)=\text{argmin}_{u^{\prime}\in\mathcal{N}(u)}\|\mathbf{x}_{u^{\prime}}-\mathbf{x}_{u_{\text{tgt}}}\|_{2}.$
@@ 4.3 Model Evaluation & Selection
To mitigate potential memorization of routes during training, especially when nodes are uniquely identifiable based on attributes, we partition the node set $V$ into three disjoint groups: $V^{\text{train}}$, $V^{\text{val}}$, and $V^{\text{test}}$ at ratios of $80\%/10\%/10\%$. The source node $u_{\text{src}}$ is sampled uniformly at random from $V$, while the target node $u_{\text{tgt}}$ depends on whether training, validation, or testing is performed. This ensures that the agent cannot memorize the path to a target node since, by construction, it is not encountered during training. We always sample a “fresh” source-target pair during training, while for validation and evaluation the source-target pairs are serialized and stored (such that the performance evaluated over them is consistent). The Mean Oracle Ratio $\bar{R}_{\text{oracle}}$ is used as the primary metric for model validation and evaluation.
@@ 5.1 Facebook Graphs
In Table 5 in the Appendix, we also include a runtime analysis to quantify the scalability of the proposed method. These results show that GARDEN maintains sub-millisecond per-action inference times even with CPU-only execution as the graph size increases, thanks to its decentralized nature. We therefore expect the method to maintain fast inference times even in substantially larger topologies.
@@ 5.3 Ablation of Node Representation
As shown in Table 2, GARDEN significantly outperforms the MLPA2C and MLPA2CWD designs that only use raw node attributes and MLP-parameterized policies in both $R_{\text{trunc}}$ and $\bar{R}_{\text{oracle}}$. The standard MLP-parameterized model achieves the best Win Rate $R_{\text{win}}$, and the differences with respect to this metric are less conclusive. This can be explained by the arbitrary tie-breaking performed when path lengths match for all methods, coupled with the high density parameter $\beta=5$ leading to shorter path lengths.
@@ A.1 Implementation Details
Our implementation is publicly available at https://github.com/flxclxc/rl-graph-search. Please see the README.md file for instructions on how to set up the dependencies, download the publicly available data, and run the code.
@@
We note that, while the GAT-based model contains 3 layers, we only use the ego graphs centered around the nodes to compute the embeddings, hence ensuring that only 1-hop visibility is provided. The first layer is fed the raw node attributes, while the second and third layers use the “latent” embeddings constructed by the previous layer. The message-passing therefore occurs over the same 1-hop subgraph in all the layers.
A.2 Additional Tables and Figures
Summary statistics. Statistics about the considered real-world graphs are shown in Table 4.
Runtime analysis. We carry out a runtime analysis to examine the scalability and computational cost of GARDEN. In terms of methodology, we calculate the mean action time, i.e., the elapsed wall clock time measured in milliseconds from the arrival of a message at a node until an action is chosen. The measurements are averaged over $100$ target nodes and are carried out using an Intel i7-11800H CPU. We note that the execution of GARDEN also involves a time overhead for creating the local ego graph embeddings from $f_{\text{rep}}^{\Theta_{3}}$, which is reported separately.
@@ A.2 Additional Tables and Figures
Graph 1 2 3 4 5
Number of Nodes 148 168 224 324 532
Avg # Neighbors 22.86 19.71 28.5 15.52 18.09
Overhead Time GARDEN (ms) 0.3024 0.2349 0.5024 0.2214 0.2143
Action Time: GARDEN (ms) 0.1592 0.1715 0.1800 0.1602 0.1656
Action Time: MLPA2CWD (ms) 0.1446 0.1548 0.1686 0.1532 0.1520
Action Time: MLPA2C (ms) 0.1499 0.1487 0.1518 0.1464 0.1469
Action Time: GreedyWalker (ms) 0.0259 0.0256 0.0323 0.0281 0.0305
Action Time: DistanceWalker (ms) 0.0657 0.0697 0.0735 0.0674 0.0685
Action Time: ConnectionWalker (ms) 0.0453 0.0522 0.0562 0.0456 0.0517
Action Time: RandomWalker (ms) 0.0006 0.0006 0.0007 0.0006 0.0006
Table 5: Runtime analysis comparing average time for calculating actions across all models, including overhead computation time of local graph embeddings for GARDEN.
The results are shown in Table 5. They demonstrate that both the embedding overhead time and action time increase slightly as the number of neighbors grows, but stays reasonably bounded. As it is expected, a clear hierarchy is present in which the neural network-based models require more inference time, followed by the heuristic attribute-based baselines, and finally the simple random walk baseline. This analysis highlights the fact that GARDEN maintains sub-millisecond inference times even with CPU-only execution as the graph size increases, thanks to its decentralized nature. Notably, the total inference time is lower on the largest 532-node graph compared to the smallest 148-node graph, owing to differences in density. We therefore expect the method to maintain fast inference times even in substantially larger topologies.
Interpretability of the learned value function. In Figure 3, we plot GARDEN’s learned value function $f_{v}^{\Theta_{2}}$ across the social network ego graphs. Brighter colors indicate a higher estimated value function relative to the target node, which is indicated with a black arrow and chosen randomly from the respective test sets. For comparison, we also plot the implicit preferability score $-\|\mathbf{x}_{u}-\mathbf{x}_{u_{\text{tgt}}}\|_{2}/\tau$ generated by the best-performing baseline, DistanceWalker, for the same source-target pairs. DistanceWalker struggles with Euclidean pairwise attribute distance due to high dimensionality and sparsity of node attributes. Conversely, values obtained by GARDEN serve as a reliable proxy for graph distance, assigning highest values to nodes in the target’s cluster or clusters with strong connectivity to the target’s community.